OpenStudy (anonymous):

The amount of money, in billions of dollars,spent on shoes for college can be estimated by polynomial 0.4t+1.13, where t is the number of years since 2004.estimate the amount spent on shoes for college in 2010

OpenStudy (anonymous):

plug in 6 for your polynomial

OpenStudy (anonymous):

t=2010-2004=6 0.4*6+1.13=3.53. Do you think this is right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Thank you so much for the help.

OpenStudy (lilg132):

yes